OpenStudy (anonymous):

A right circular cone has a height of 3.75 in. The diameter of the base is 1.2 in. Find the volume of the cone.

jhonyy9 (jhonyy9):

V=area of base *h/3 D=2r r=D/2

OpenStudy (anonymous):

WHAT?

jhonyy9 (jhonyy9):

r=D/2 =1,2/2 = 0,6 Area of base = pir2 = pi(0,6*0,6)=0,36*pi V=(0,36pi*3,75)/3 = 0,45pi in squared

jhonyy9 (jhonyy9):

sorry 0,45 in cubed